Sqrt x^3=27
Explain the answer
^3 sqrt x^2=9
Show work below

Answer by stanbon(75887)

Sqrt x^3=27
Square both sides to get rid of the radical:
x^3=27^2
Take the cube root to get:
x=3^2
x=9
----------------------
Explain the answer
^3 sqrt x^2=9
Comment: What do you mean by ^3 at the beginning of the line?
Show work below
---------
Cheers,
Stan H.